# label*=0., itemsep=5pt, topsep=5pt

### v1

@@ Abstract.
The paper concerns the coarse flag Hilbert-Poincaré series of [4] in the case of the braid arrangement $\mathcal{B}_{n}$ associated to the symmetric group ${\mathfrak{S}}_{n+1}$. We explicitly construct a companion statistic $\operatorname{ino}:{\mathfrak{S}}_{n+1}\times\operatorname{Sym}(n)\rightarrow\mathbb{N}$ for the descent statistic on $\operatorname{Sym}(n)$ using reverse $(P,\omega)$-partitions and quasisymmetric functions.
@@ 1.1. Background
where we set $C_{0}=V$ and $C_{k+1}=\mathbf{0}$. The coarse flag Hilbert-Poincaré series is defined by Maglione–Voll in [4] as
@@
where the sum ranges over all chains in the intersection poset $\mathcal{L}$ not containing the bottom element $V$. Its numerator polynomial has nonnegative coefficients, $\mathcal{N}_{\mathcal{A}}(y,t)\in\mathbb{N}[y,t]$, see [2]. It is moreover shown in [3] for simplicial arrangements that
@@
If $\mathcal{A}$ is the reflection arrangement of a real reflection group $W$, the regions are in bijection with the elements of the group $W$, so this becomes the question of finding a companion statistic on $W\times\operatorname{Sym}(n)$. The purpose of this paper is to answers this question for the symmetric group. The given companion statistic moreover provides a new approach towards generalizing left-to-right minima on permutations to real reflection groups, see Remark 1.5.
[…]
@@ 1.2. Main result
The merging blocks $3|1$ in the cover relation $3|1|2\prec 13|2$ are not in order, while the merging blocks in $13|2\prec 123$ are. Thus, $\operatorname{ino}({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}312},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}12})=1$. All the blocks in the blue chain $({\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}123},{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}21})$ come in order, so $\operatorname{ino}({\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}123},{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}21})=2$.
@@
For a real reflection group $W$, 1.1 becomes the question of providing a companions statistic for descents on $\operatorname{Sym}(n)$ in terms of the group $W$,
@@
It is known that $\mathcal{N}_{\mathcal{A}}(y,0)=\operatorname{Poin}(\mathcal{A},y)$, where the latter is the Poincaré polynomial of the arrangement, see [4]. For reflection arrangements $\mathcal{A}_{W}$, this corresponds to
@@ 1.3. Refining the main result
In order to prove Theorem 1.4, we make use of a formula for the numerator polynomial $\mathcal{N}_{\mathcal{A}}(y,t)$ based on an $R$-labeling (see Section 2.1) on the lattice of flats of $\mathcal{B}_{n}$. The set partition lattice $\Pi_{n}$ has an $R$-labeling $\lambda$ given by labeling a cover relation merging two blocks $B$ and $B^{\prime}$ by $\max\big{\{}\min(B),\min(B^{\prime})\big{\}}$, see [1, Example 2.9]. We extend this $R$-labeling from the set partition lattice to an edge-labeling of the poset of set compositions $\Sigma_{n}$ by defining $\lambda(w,\sigma)=(\lambda_{1},\dots,\lambda_{n})$ for $w\in{\mathfrak{S}}_{n+1}$ and $\sigma\in\operatorname{Sym}(n)$ by
@@
(Here and below, we often write $\overline{x}=-x$ for $x\in\mathbb{Z}$.) The sequences $\lambda(w,\sigma)$ for the chains $({\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}123},{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}21})$ and $({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}312},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}12})$ are in particular
[…]
We obtain $\operatorname{asc}\big{(}\lambda({\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}123},{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}21})\big{)}=0$ and $\operatorname{asc}\big{(}\lambda({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}231},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}12})\big{)}=1$.
@@ 2. Proofs
  3. (3)
    collect two properties of the $R$-labeling (Section 2.3),
  4. (4)
@@ 2.1. Rewriting the numerator polynomial
Let $P$ be a finite and ranked poset of rank $n$ with unique minimal element $\widehat{0}$ and unique maximal element $\widehat{1}$. Denote its set of cover relations by $E(P)=\left\{(x,y)\;|\;x\prec y\right\}$. An edge-labeling $\lambda:E(P)\to\mathbb{N}_{+}$ is an $\boldsymbol{R}$-labeling if every interval in $P$ contains a unique $\lambda$-weakly increasing maximal chain. Throughout this section, we assume the edge-labeling to be injective along maximal chains. For a maximal chain =$\mathcal{M}=\{M_{0}\prec\dots\prec M_{n}\}$, define its labeling as $\lambda(\mathcal{M})=\big{(}\lambda(\mathcal{M}_{0},\mathcal{M}_{1}),\dots,\lambda(\mathcal{M}_{n-1},\mathcal{M}_{n})\big{)}$. For a subset $Y\subseteq\{1,\dots,n\}$, we slightly modify $\lambda(\mathcal{M})$ to obtain $\lambda(\mathcal{M},Y)=\big{(}\!\pm\lambda(\mathcal{M}_{0},\mathcal{M}_{1}),\dots,\pm\lambda(\mathcal{M}_{n-1},\mathcal{M}_{n})\big{)}$ where the sign is positive if $i\in Y$ and negative if $i\notin Y$.
The lattice of flats $\mathcal{L}$ of a hyperplane arrangement $\mathcal{A}$ always admits an $R$-labeling $\lambda$ [5, Example 3.14.5]. The numerator polynomial $\mathcal{N}_{\mathcal{A}}$ was shown in [2] to have the description
@@
where the sum ranges over all maximal chain $\mathcal{M}$ in $\mathcal{L}$ and all subsets $Y\subseteq\{1,\dots,n\}$ for $n=\operatorname{rk}(\mathcal{L})$, and where $\operatorname{asc}\big{(}\lambda(\mathcal{M},Y)\big{)}$ denotes the number of ascents of $\lambda(\mathcal{M},Y)$. We refer to [2] for further details.
@@
2.2. Further refining the main result
@@ 2.3. Behavior of the edge-labeling $\boldsymbol{\lambda}$
Throughout this section, fix $w\in{\mathfrak{S}}_{n+1}$ and $\sigma\in\operatorname{Sym}(n)$. Moreover, recall the definition of $\lambda(w,\sigma)$ given in (1.1).
@@
$\lambda(w,\sigma)=($ $\overline{2},$ $6,$ $5,$ $\overline{4},$ $3$ ).))\,.) .
@@ 2.5. Quasisymmetric functions and a proof of the main result
In order to prove Theorem 2.2, we use the following implication of the theory of reverse $(P,\omega)$-partitions from [6, §7.19].
@@
We are left to provide a proof of Proposition 2.11. Let $P$ be a finite poset on $\{1,\dots,n\}$ and let $\omega:P\to\mathbb{Z}$ be an injective vertex-labeling of $P$. Following [6, §7.19], a reverse $\boldsymbol{(P,\omega)}$-partition is a map $\sigma:P\to\mathbb{N}$ satisfying
  1. (1)

### v2

# bibliography.bib

@@ Abstract.
[…]
@@ 1.1. Background
where we set $C_{0}=V$ and $C_{k+1}=\mathbf{0}$. The coarse flag Hilbert-Poincaré series is defined by Maglione–Voll in [MaglioneVoll] as
@@
where the sum ranges over all chains in the intersection poset $\mathcal{L}$ not containing the bottom element $V$. Its numerator polynomial has nonnegative coefficients [poincareextended] and the Chow polynomial of the associated matroid is given by the evaluation $\operatorname{cfHP}_{\mathcal{A}}(-x,x)$ [stump2024chowaugmentedchowpolynomials]. It is moreover shown in [KühneMaglione] for simplicial arrangements that
@@
If $\mathcal{A}$ is the reflection arrangement of a real reflection group $W$, the regions are in bijection with the elements of the group $W$, so this becomes the question of finding a companion statistic on ${W\times\operatorname{Sym}(n)}$. We emphasize that it already appears to be a challenging task to find such a companion statistic, and that this statistic would in particular provide a new approach towards generalizing left-to-right minima on permutations to real reflection groups, see Remark 1.5.
The purpose of this paper is to introduce a new statistic on permutations which answers 1.1 for the symmetric group.
Let $\mathcal{B}_{n}$ be the braid arrangement given by the hyperplanes $H_{ij}=\{x_{i}=x_{j}\}$ for $11\leq i<j\leq n+$ 1 in $V=\big{\{}(x_{1},\dots,x_{n+1})\in\mathbb{R}^{n+1}\mid x_{1}+\dots+x_{n+1}=0\big{\}}\cong\mathbb{R}^{n}$, the corresponding reflection group is $W={\mathfrak{S}}_{n+1}$. Its lattice of flats, denoted by $\Pi_{n}$, is given by the lattice of set partitions of $\{1,\dots,n+1\}$ of rank $n$, with cover relations being given by merging two blocks. Its minimal element is the all-singletons set partition and its maximal element is the all-in-one-block set partition. A set composition is a set partition together with a linear ordering of the blocks. We write $B_{1}|B_{2}|\cdots|B_{k}$ for the set composition with the blocks being ordered from left to right. The face poset of the braid arrangement $\mathcal{B}_{n}$, denoted by $\Sigma_{n}$, is given by the poset of set compositions of $\{1,\dots,n+1\}$ of rank $n$, with cover relations being given by merging two consecutive blocks. Its minimal elements are thus in natural bijection with permutations, and its unique maximal element is the all-in-one-block set composition.
@@ 1.2. Main result
The merging blocks $3|1$ in the cover relation $3|1|2\prec 13|2$ are not in order, while the merging blocks in $13|2\prec 123$ are. Thus, $\operatorname{ino}({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}312},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}12})=1$. All the blocks in the chain $({\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}123},{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}21})$ come in order, so $\operatorname{ino}({\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}123},{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}21})=2$.
@@
For a real reflection group $W$, 1.1 becomes the question of providing a companion statistic for descents on $\operatorname{Sym}(n)$ in terms of the group $W$,
@@
It is known that $\mathcal{N}_{\mathcal{A}}(y,0)=\operatorname{Poin}(\mathcal{A},y)$, where the latter is the Poincaré polynomial of the arrangement, see [MaglioneVoll]. For reflection arrangements $\mathcal{A}_{W}$, this corresponds to
@@ 1.3. Refining the main result
In order to prove Theorem 1.4, we make use of a formula for the numerator polynomial $\mathcal{N}_{\mathcal{A}}(y,t)$ based on an $R$-labeling (see Section 2.1) on the lattice of flats of $\mathcal{B}_{n}$. The set partition lattice $\Pi_{n}$ has an $R$-labeling $\lambda$ given by labeling a cover relation merging two blocks $B$ and $B^{\prime}$ by ````, see [MR0570784, Example 2.9]. We extend this $R$-labeling from the set partition lattice to an edge-labeling of the poset of set compositions $\Sigma_{n}$ by defining $\lambda(w,\sigma)=(\lambda_{1},\dots,\lambda_{n})$ for $w\in{\mathfrak{S}}_{n+1}$ and $\sigma\in\operatorname{Sym}(n)$ by
@@
(Here and below, we often write $\overline{x}=-x$ for $x\in\mathbb{Z}$.) The sequences $\lambda(w,\sigma)$ for the chains $({\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}123},{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}21})$ and $({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}312},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}12})$ are in particular
$\lambda({\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}123},{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}21})=(3,2)\quad\text{and}\quad\lambda({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}312},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}12})=(\overline{3},2)\,.$
We obtain $\operatorname{asc}\big{(}\lambda({\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}123},{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}21})\big{)}=0$ and $\operatorname{asc}\big{(}\lambda({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}231},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}12})\big{)}=1$.
@@
Remark 1.10.
As written in [MR0570784, Example 2.9], the “max-of-min” labeling in (1.1) is the edge-labeling given in [MR0354473, Proposition 2.4] for supersolvable lattices, using the $\mathcal{M}$-chain
[…]
in $\Pi_{n}$. This $\mathcal{M}$-chain induces the admissible map $H_{ij}\mapsto j$, which in turn induces the “max-of-min” $R$-labeling. This admissible map, and thus the “max-of-min” $R$-labeling, may as well be obtained by considering the induced edge-labeling of an admissible map given in [MR0354473, Proposition 2.2] for geometric lattices, using the total order
$H_{12}<H_{13}<H_{23}<H_{14}<H_{24}<H_{34}<\dots<H_{1(n+1)}<\dots<H_{n(n+1)}$
of the atoms in $\Pi_{n}$. It is easy to check that no two labels $H_{ij}$ and $H_{i^{\prime}j}$ can occur along a maximal chain in $\Pi_{n}$, implying again that $H_{ij}\mapsto j$ is an admissible map.
@@ 2. Proofs
  3. 3.
    collect several properties of the $R$-labeling and reestablish a symmetry of $\mathcal{N}_{\mathcal{B}_{n}}$ (Section 2.3),
  4. 4.
@@ 2.1. Rewriting the numerator polynomial
Let $P$ be a finite and ranked poset of rank $n$ with unique minimal element $\widehat{0}$ and unique maximal element $\widehat{1}$. Denote its set of cover relations by $E(P)=\left\{(x,y)\;|\;x\prec y\right\}$. An edge-labeling $\lambda:E(P)\to\mathbb{N}_{+}$ is an $\boldsymbol{R}$-labeling if every interval in $P$ contains a unique $\lambda$-weakly increasing maximal chain. Throughout this section, we assume the edge-labeling to be injective along maximal chains. For a maximal chain =$\mathcal{M}=\{\mathcal{M}_{0}\prec\dots\prec\mathcal{M}_{n}\}$, define its labeling as $\lambda(\mathcal{M})=\big{(}\lambda(\mathcal{M}_{0},\mathcal{M}_{1}),\dots,\lambda(\mathcal{M}_{n-1},\mathcal{M}_{n})\big{)}$. For a subset $Y\subseteq\{1,\dots,n\}$, we slightly modify $\lambda(\mathcal{M})$ to obtain $\lambda(\mathcal{M},Y)=\big{(}\!\pm\lambda(\mathcal{M}_{0},\mathcal{M}_{1}),\dots,\pm\lambda(\mathcal{M}_{n-1},\mathcal{M}_{n})\big{)}$ where the sign is positive if $i\in Y$ and negative if $i\notin Y$.
The lattice of flats $\mathcal{L}$ of a hyperplane arrangement $\mathcal{A}$ always admits an $R$-labeling $\lambda$ [MR2868112, Example 3.14.5]. The numerator polynomial $\mathcal{N}_{\mathcal{A}}$ was shown in [poincareextended] to have the description
@@
where the sum ranges over all maximal chain $\mathcal{M}$ in $\mathcal{L}$ and all subsets $Y\subseteq\{1,\dots,n\}$ for $n=\operatorname{rk}(\mathcal{L})$, and where $\operatorname{asc}\big{(}\lambda(\mathcal{M},Y)\big{)}$ denotes the number of ascents of $\lambda(\mathcal{M},Y)$. We refer to [poincareextended] for further details.
@@
2.2. Further refinement of the main result
@@ 2.3. Behavior of the edge-labeling $\boldsymbol{\lambda}$
Throughout this section, fix $w=w_{1}\dots w_{n+1}\in{\mathfrak{S}}_{n+1}$ and $\sigma\in\operatorname{Sym}(n)$. Moreover, recall the definition of $\lambda(w,\sigma)$ given in (1.1). We start with the following two lemmas, for which we set $\overleftarrow{w}=w_{n+1}\cdots w_{1}\in{\mathfrak{S}}_{n+1}$ sending $i$ to $w_{n+2-i}$ and ${\sigma_{\uparrow}\in\operatorname{Sym}(n)}$ sending $i$ to $n+1-\sigma(i)$.
Lemma 2.4.
Let $\lambda(w,\sigma)=(\lambda_{1},\dots,\lambda_{n})$. Then
$\lambda(\overleftarrow{w},\sigma_{\uparrow})=(-{\lambda_{1}},\dots,-{\lambda_{n}})\,.$
Proof.
The $i$-th cover relation in $(w,\sigma)$ merges the exact same blocks as the $i$-th cover relation in $(\overleftarrow{w},\sigma_{\uparrow})$. The merging blocks only come in opposite order, swapping the sign of $\lambda_{i}$. ∎
@@
Combining these two lemmas yields the following symmetry on ¸$\mathcal{N}_{\mathcal{B}_{n}}$. More general forms of this symmetry have already been observed in  [MaglioneVoll, Corollary 3.3] and [poincareextended, Corollary 2.16].
Corollary 2.6.
It holds that
$\operatorname{Ino}(\overleftarrow{w},\sigma_{\uparrow})$ =$=\{2,\dots,n+1\}\setminus\operatorname{Ino}(w,\sigma)$
$\operatorname{Asc}\big{(}\lambda(\overleftarrow{w},\sigma_{\uparrow})\big{)}$ =$=\{2,\dots,n+1\}\setminus\operatorname{Asc}\big{(}\lambda(w,\sigma)\big{)}\,.$
@@
$\lambda(w,\sigma)=($$2$ $\overline{2},$ $6,$ $5,$ $\overline{4},$ $3$ ).))\,.) .$123456\,.$
@@ 2.5. Quasisymmetric functions and a proof of the main result
In order to prove Theorem 2.2, we use the following implication of the theory of reverse $(P,\omega)$-partitions from [MR4621625, §7.19].
@@
We are left to provide a proof of Proposition 2.13. Let $P$ be a finite poset on $\{1,\dots,n\}$ and let $\omega:P\to\mathbb{Z}$ be an injective vertex-labeling of $P$. Following [MR4621625, §7.19], a reverse $\boldsymbol{(P,\omega)}$-partition is a map $\sigma:P\to\mathbb{N}$ satisfying
  1. 1.